# A Guide to Misinformation Detection Datasets

### v1

@@ A.2 Dataset Details
CMU-MisCOV19 (Memon & Carley, 2020): CMU-MisCov19 is a dataset about COVID-19. It contains tweets that were collected over three days: March 28, 2020, June 15, 2020, and June 24, 2020. 4,573 tweets are annotated based on various types of information and misinformation. In total, there are 17 categories, such as irrelevant, conspiracy, true treatment, fake cure, false fact, ambiguous, etc.
CoAID (Cui & Lee, 2020): This dataset covers various COVID-19 healthcare misinformation. It contains 4,251 news, 926 social platforms posts, and 296,000 related user engagements. All facts are collected between December 1, 2019 and September 1, 2020. All the data is annotated in a binary form: true or fake.
@@
FakeCovid (Shahi & Nandini, 2020): FakeCovid is a dataset containing news claims about COVID-19. These data were collected from 92 different fact-cheking websites between January 4, 2020, and May 15, 2020, covering 40 languages and originating from 105 countries. The truthfulness labels (false, mostly false, misleading, half true, mostly true, no evidence) are derived from experts at fact-checking agencies. The dataset also includes other labels defining the type of false news (prevention & treatments, international response, conspiracy theories, etc), all annotated by members of their team.
@@
IFND (Sharma & Garg, 2023): The Indian Fake News Dataset (IFND) consists of texts and images collected between 2013 and 2021. These data cover elections, politics, COVID-19, violence, and miscellaneous topics. The veracity of these data is determined based on the media from which they were collected. True claims originate from Tribune, Times Now News, The Statesman, and others, while false claims come from the fact-checked columns of Alt News, Boomlive, and media outlets like The Logical Indian, and News Mobile.
@@
X-Fact (Gupta & Srikumar, 2021): X-FACT is a dataset of 31,189 short statements scraped from 85 fact-checking websites. Covering various topics, the data are available in 25 languages, including Arabic, Bengali, French, Hindi, Indonesian, Italian, Spanish, Polish, and Portuguese. The veracity labels indicate a decreasing level of truthfulness: true, mostly true, partly true, mostly false, false, unverifiable, and other.
A.3 Supplement on Keyword Analysis
@@
Figures  2,  4,  6 and  8 show the distribution of the 40 most frequent words across the IFND, MM-COVID, Truthseeker2023, and Twitter16 datasets. The prevalence of the words in each veracity category was calculated using their relative frequency. A word positioned at x = 1 indicates that it is systematically associated with the veracity category specified by the label, meaning that 100% of statements containing this word are labeled the same way. Additionally, figures  24,  6, and  8, plotted using ScatterText, highlight the words that had a significant impact on the random classifier presented in Table  3, with color indicating the frequency of a word’s association with a label.
Table 7: Identification of spuriously predictive keywords.
@@
Figure 1: IFND
Figure 2: IFND Predictivity
Figure 3: MM-COVID
Figure 4: MM-COVID Predictivity
Figure 5: Truthseeker2023
Figure 6: Truthseeker2023 Predictivity
Figure 7: Twitter16
Figure 8: Twitter16 Predictivity
A.4 Supplement on Political Leaning
The meaning of ideology and its dimensions varies across countries and regions Caughey et al. (2019); Bauer et al. (2017); Lo et al. (2014). However, most datasets, including the most popular ones on claims, focus on statements made in the U.S. or pertaining to the American context (see Table 1). Given this and the polarization of the American political landscape, we use a partisanship scale that classifies statements as either leaning Republican, leaning Democratic, leaning neither Republican nor Democratic or not political. This allows us to identify the proportion of true and false claims in each dataset that fall within these four categories. More specifically, we use the following prompt on gpt-4-turbo-0409:
Statement: {{Statement to analyze}}
Your task is to analyze the partisan lean of different statements. Think step-by-step. First, indicate whether the statement is more favorable to Republicans, more favorable to Democrats, neither more favorable to Republicans or Democrats or not political. For instance, if a statement supports a position commonly associated with the Republican Party, it might be more favorable to Republicans. Similarly, if a statement highlights a negative characteristic of a Democratic candidate/political figure or a positive characteristic of a Republican candidate/political figure it might also be more favorable to Republicans.
After providing all of your analysis steps, summarize your analysis and state: "Republican Lean; Partisan Lean: (A)" if you think the statement is more favorable to Republicans, or "Democratic Lean; Partisan Lean: (B)" if you think the statement is more favorable to Democrats, or "Neither Republican or Democratic Lean; Partisan Lean: (C)" if you think the political statement is not more favorable to Republicans or Democrats, or "Non Political; Partisan Lean: (D)" if you think the statement is not political. You should begin your summary with the phrase "Summary:"
Analyzing the partisan lean of statements is crucial for several reasons. First, providing distributions of the partisan lean enables researchers to select datasets that better align with their specific research goals. For instance, researchers interested in analyzing a balance of Democrat—and Republican—leaning misinformation can choose datasets accordingly. This targeted selection improves the relevance and precision of their studies. Second, if the veracity of a statement correlates with its partisan lean, models may inaccurately assess the statement’s veracity based solely on that lean.
Table 8 summarizes our findings. Overall, we observe that a slightly higher proportion of true statements across the datasets are predicted to lean Democratic (12.12%) compared to those leaning Republican (9.5%). In contrast, a greater proportion of false statements are predicted to lean Republican (19.44%) than Democratic (9.36%). However, these proportions vary significantly across datasets. For instance, this partisan bias appears to be more pronounced in some of the most commonly used political datasets, such as LIAR and Twitter 15/16. However, in some datasets—particularly those related to vaccines and COVID-19 (e.g., AntiVax, Check-COVID, MM-COVID)—a larger proportion of statements are classified as neither Democratic nor Republican leaning or as non-political. This is understandable, as anti-vaccine attitudes often emerge from both sides of the political spectrum, especially in the earlier phase of the pandemic (Roberts et al., 2022). Finally, we note that statements that are not explicitly related to politics tend to be more true on average (57.51% vs. 48.28%). This suggests a potential bias on models trained with these data, where political statements may be assumed to be more likely false.
Finally, we note our measurement approach does not account for the context surrounding a claim, it offers a prediction of the partisan lean based solely on the claim itself. Without context, subtleties such as veracity, sarcasm, or nuance might be missed, potentially affecting the predictions. Future research should explore how incorporating additional context might enhance the accuracy and clarity of these predictions. One approach could be to use retrieval-augmented generation by integrating external information from reliable sources for context (Kortukov et al., 2024). Without this added information, models could make incorrect predictions because of the limited context of certain claims. It is thus crucial not only to consider the distribution of the partisan lean when using datasets, but also to recognize that predictive models may lack information on the broader political context when evaluating the veracity of claims.
Table 8: True and False Statements Percentages by Political Leanings
Dataset % Rep. Lean (True) % Dem. Lean (True) % Neither (True) % Not Political (True) % Rep. Lean (False) % Dem. Lean (False) % Neither (False) % Not Political (False)
IFND 9.1 10.75 34.78 45.37 9.63 6.21 42.55 41.61
AntiVax 0 0 0.26 99.74 0.16 0 0 99.84
Check-COVID 4.27 13.87 41.33 40.53 23.71 4.63 38.15 33.51
Climate-Fever 21.22 41.49 20.08 17.21 67.5 8.12 11.88 12.5
CMU-MisCOV19 0 0 0 100 0 0 0.32 99.68
CoAID 3.1 8.34 44.49 44.06 52.38 4.76 28.57 14.29
Counter-covid-19-misinformation 0 0 0 0 0 0 0 100
Covid-19-disinformation 0 0 0.14 99.86 0 0 0 100
COVID-19-Rumor 13.75 11.15 32.34 42.75 17.45 7.48 39.2 35.87
COVID-Fact 3.25 6.49 29.55 60.71 5.54 5.54 34.11 54.81
ESOC Covid-19 0 0 0 0 12.12 10.08 42.06 35.74
FaVIQ 0.27 1.49 13.15 85.08 0.8 0.8 14.32 84.08
FEVER 1.29 1.47 14 83.24 0.94 0 16.51 82.55
FEVEROUS 0.84 1.87 11.87 85.42 2.05 1.57 14.72 81.66
FibVID 31.67 38.33 20.83 9.17 49.93 20.53 18.01 11.52
HoVer 0.46 1.47 12.72 85.35 0.67 0.89 11.74 86.69
LIAR 37.72 31.53 24.82 5.93 51.38 25.69 18.25 4.68
LIAR-New 44.68 34.04 19.15 2.13 53.24 16.14 19.2 11.43
MediaEval 0.22 0 0.22 99.56 0 0 0.39 99.61
MM-COVID 1.27 8.2 13.15 77.37 22.22 9.32 39.43 29.03
PHEME 17.37 22.14 32.3 28.19 10.23 18.6 37.67 33.49
PubHealthTab 4.9 6.15 31.47 57.48 4.44 5.92 31.95 57.69
Rumors 18.92 22.52 23.42 35.14 25.96 15.68 27 31.36
Snopes Fact-news 21.08 22.89 28.31 27.71 28.81 15.58 30.32 25.29
TruthSeeker2023 26.37 42.97 25 5.66 72.82 8.09 14.32 4.77
Twitter15 3.21 4.02 15.26 77.51 13.83 8.5 18.18 59.49
Twitter16 4.83 17.87 23.19 54.11 20.25 12.84 17.04 49.88
Verite 6.53 13.65 30.56 49.26 14.95 11.93 32.78 40.33
WICO 0 0 0.31 99.69 0 0 0.3 99.7
X-Fact 11.72 15.45 35.85 36.98 15.86 13.32 35.33 35.49
Overall 9.5 12.12 20.87 57.51 19.44 9.36 22.93 48.28
A.5 Implementation Details of GPT-4 with Web Search Predictive System

### v2

@@ A.7.1. Claims Dataset Details
CMU-MisCOV19 (Memon and Carley, 2020): CMU-MisCov19 is a dataset about COVID-19. It contains tweets that were collected over three days: March 28, 2020, June 15, 2020, and June 24, 2020. 4,573 tweets are annotated based on various types of information and misinformation. In total, there are 17 categories, such as irrelevant, conspiracy, true treatment, fake cure, false fact, ambiguous, etc.
CoAID (Cui and Lee, 2020): This dataset covers various COVID-19 healthcare misinformation. It contains 4,251 news, 926 social platforms posts, and 296,000 related user engagements. All facts are collected between December 1, 2019 and September 1, 2020. All the data is annotated in a binary form: true or fake.
@@
FakeCovid (Shahi and Nandini, 2020): FakeCovid is a dataset containing news claims about COVID-19. These data were collected from 92 different fact-cheking websites between January 4, 2020, and May 15, 2020, covering 40 languages and originating from 105 countries. The truthfulness labels (false, mostly false, misleading, half true, mostly true, no evidence) are derived from experts at fact-checking agencies. The dataset also includes other labels defining the type of false news (prevention & treatments, international response, conspiracy theories, etc), all annotated by members of their team.
@@
IFND (Sharma and Garg, 2023): The Indian Fake News Dataset (IFND) consists of texts and images collected between 2013 and 2021. These data cover elections, politics, COVID-19, violence, and miscellaneous topics. The veracity of these data is determined based on the media from which they were collected. True claims originate from Tribune, Times Now News, The Statesman, and others, while false claims come from the fact-checked columns of Alt News, Boomlive, and media outlets like The Logical Indian, and News Mobile.
@@
X-Fact (Gupta and Srikumar, 2021): X-FACT is a dataset of 31,189 short statements scraped from 85 fact-checking websites. Covering various topics, the data are available in 25 languages, including Arabic, Bengali, French, Hindi, Indonesian, Italian, Spanish, Polish, and Portuguese. The veracity labels indicate a decreasing level of truthfulness: true, mostly true, partly true, mostly false, false, unverifiable, and other.
A.7.2. Paragraph Dataset Details
BanFakeNews. (Hossain et al., 2020): BanFakeNews is a dataset containing approximately 50,000 news articles. The articles are in the Bangla language and are collected from 22 popular news portals in Bangladesh. The articles were labelled by crowd-sourcing. This dataset is annotated with a set of labels consisting of fake and authentic.
BenjaminPoliticalNews. (Horne and Adali, 2017): BenjaminPoliticalNews is a dataset containing 296 news articles. These articles were collected from existing datasets and studies. The articles were labelled by a source-based method. This dataset is annotated with a set of labels consisting of real, satire, fake, and true.
Celebrity. (Pérez-Rosas et al., 2017): Celebrity is a dataset containing 500 news articles from magazines about celebrity gossip and hoaxes. The articles were labelled by a source-based method. This dataset is annotated with a set of labels consisting of legit and fake.
CT-FAN. (Shahi et al., 2021): CT-FAN is a dataset containing 2462 news articles. These articles were collected from multiple fact-checking sites with news articles from 2010 to 2022. The articles were labelled by a source-based method. This dataset is annotated with a set of labels consisting of partially false, false, other, and true.
FA-KES. (Abu Salem et al., 2019): FA-KES is a dataset containing 804 news articles. These articles were collected from multiple fact-checking websites and social media platforms, encompassing multiple languages and domains. The articles were labelled by crowd-sourcing. This dataset is annotated with a set of labels consisting of true, authentic, and fake.
FakeNewsCorpus. (Pathak and Srihari, 2019): FakeNewsCorpus is a dataset containing 9408908 articles. These articles were collected from a curated list of 1001 domains. The articles were labelled by a source-based method. This dataset is annotated with a set of labels consisting of unreliable, fake, clickbait, conspiracy, reliable, bias, hate, junksci, political, unknown, and nan.
FakeNewsAMT. (Pérez-Rosas et al., 2017): fakenewsamt is a dataset containing 480 news articles. Real news articles were sourced from reputable news websites, while fake news articles were generated using Amazon Mechanical Turk, where crowdworkers were tasked with writing fictitious news content on given topics. This dataset is annotated with a set of labels consisting of fake and true.
ISOT Fake News Dataset. (of Victoria, 2022): ISOT Fake News Dataset is a dataset containing 44898 news articles. Real news articles were collected by crawling Reuters.com, while fake news articles were collected from unreliable websites identified by PolitiFact.com. This dataset is annotated with a set of labels consisting of true and false.
TI-CNN. (Yang et al., 2023): TI-CNN is a dataset containing 20015 news articles. These articles were collected from social media platforms and news websites. The articles were labelled by a source-based method. This dataset is annotated with a set of labels consisting of fake and real.
A.8. Supplement on Labeling Approach
Table 5. Labeling approach and label distribution for 36 claim datasets and 9 paragraph datasets (subset of Table 4).
Dataset Labeling approach True (%) False (%) Mixed (%) Unknown (%)
AntiVax Human expert 38.15 61.85 - -
Check-COVID Human expert 37.92 37.16 - 24.92
ClaimsKG Human expert 17.23 63.06 12.32 7.39
Climate-Fever Human expert 42.61 16.48 10.03 30.88
CMU-MisCOV19 Human (N.S.) 7.39 70.63 - 21.98
CoAID Source-based categorization (T) & Human expert (F) 93.47 6.53 - -
Counter-covid-19-misinformation Human (N.S.) 1.08 1.09 - 97.83
COVID-19-Rumor Human expert 26.16 51.27 - 22.57
Covid-19-disinformation Crowd-sourced - - - 100
COVID-Fact Human-expert (T) & Algorithm-generated creation (F) 31.72 68.28 - -
Covid-vaccine-misinfo-MIC Crowd-sourced - - - 100
DeFaktS Human expert 11.12 7.78 - 81.1
ESOC Covid-19 Human expert - 99.52 - 0.48
FakeCovid Human expert 0.85 94.12 2.74 2.28
FaVIQ Algorithm & Validation by human 49.77 50.23 - -
FEVER Crowd-sourced 46.75 19.65 - 33.6
FEVEROUS Algorithm 57.77 38.77 - 3.47
FibVID Human expert 23.76 76.24 - -
HoVer Crowd-sourced 49.76 34.95 - 15.28
IFND Human expert 66.64 33.36 - -
LIAR Human expert 52.29 27.95 19.7 0.06
LIAR-New Human expert 19.62 72.87 7.51
MediaEval Source-based categorization 44.31 51.92 - 3.77
MM-COVID Human expert 71.74 28.26 - -
MultiClaim Human expert - 57.66 16.49 25.85
NLP4IF-2021 Crowd-sourced 64.31 3.28 - 32.41
PHEME Human expert and non-expert 33.77 66.23 - -
PubHealthTab Human expert 52.47 23.79 - 23.74
Rumors Algorithm 11.25 58.25 6.47 24.03
Snopes Fact-news Human expert 16.07 65.41 10.95 7.58
TruthSeeker2023 Crowd-sourced 51.36 48.64 - -
Twitter15 Human expert 50.07 24.83 - 25.1
Twitter16 Human expert 50.37 25.06 - 24.57
Verite Human expert 33.77 66.23 - -
WICO Human expert 68.32 31.68 - -
X-Fact Human expert 30.26 59.64 6.33 3.78
BanFakeNews crowd-sourced 95.56 4.44 - -
Benjamin Political News Dataset source-based 39.26 60.74 - -
Celebrity source-based 50.00 50.00 - -
CT-FAN source-based 26.97 64.78 - 8.25
FA-KES crowd-sourced 52.99 47.01 - -
FakeNewsCorpus source-based 1.20 79.20 - 19.60
FakeNewsAMT crowd-sourced 50.00 50.00 - -
ISOT Fake News source-based 47.70 52.30 - -
TI-CNN source-based 40.34 59.66 - -
(T) indicates the method used to establish true claims
(F) indicates the method used to determine false claims
(N.S.) indicates that expertise is not specified
The task of annotating statements is both crucial and challenging for anyone attempting to train a robust classifier for misinformation detection. Precise labeling is essential to ensure the classifier’s effectiveness, as it directly impacts its performance and reliability. Numerous approaches have been proposed in the literature to label true and false information. These approaches include expert and crowd-sourced annotation, source-based techniques, algorithmic methods, and a hybrid of these different approaches, all of which have been used in at least one of our 36 datasets (see Table 5). We thus describe these different approaches used by the authors of the original datasets in turn to highlight their potential advantages and limitations.
Expert-based approach
Experts and fact-checkers are a small group of non-partisan professionals from various disciplines who manually verify the veracity of information. The result of these verifications are often published in fact-checking websites such as Politifact or Snopes. The strength of this approach lies in its rigorous review process, ensuring each piece of information is thoroughly evaluated, which leads to consistent reviews across fact-checkers. However, this method is not scalable and is costly (Zhou and Zafarani, 2020). As a result, experts must selectively choose the information they evaluate, which leads to many pieces of information going unchecked and potential biases in the selection of news and information that is evaluated (Lee et al., 2023; Markowitz et al., 2023; Walker and Gottfried, 2019).
Crowd-sourced approach
Crowdsourced fact-checking involves enlisting non-expert laypeople to assess the accuracy of online information. These evaluations are then aggregated to determine the veracity of the content. This approach is advantageous because it is more scalable, and laypeople can respond to misinformation much more quickly than professional fact-checkers (Zhao and Naaman, 2023). Additionally, this method has been shown to be effective in reducing the spread of misinformation and to produce veracity ratings similar to those of professional fact-checkers (Allen et al., 2021; Martel et al., 2024). However, crowdsourcing also has its limitations. It can be challenging to filter out evaluations from non-credible users and to ensure a balanced representation of users from different partisan backgrounds (Zhou and Zafarani, 2020; Martel et al., 2024).
Source-based approach
Source-based approaches to verifying information involve evaluating the domain or author of the content. Information is then rated as accurate if it comes from reliable sources and inaccurate otherwise. This method is more scalable than manual fact-checking, as it consists of evaluating the credibility of the source rather than each individual story. Additionally, this method is proven to be reliable, as experts generally rate news domains similarly (Lin et al., 2023). However, there are notable drawbacks. For instance, individual stories can vary in accuracy even within the same source, and not all content from low-quality outlets is necessarily false or misleading. Additionally, source familiarity significantly influences the perceived trustworthiness of content. Sources that are unfamiliar are often less trusted, which can lead to unfair negative evaluations of high-quality but lesser-known sources (Pennycook and Rand, 2019; Williams-Ceci et al., 2023)
Algorithmic methods
Finally, algorithmic methods can also be used to evaluate the veracity of content using NLP or other ML techniques (Zhou and Zafarani, 2020). For example, Covid-fact uses a BERT-based classifier, FaVIQ uses T5-3B, and Rumors uses an approach based on a social graph. These methods offer significant advantages in scalability, as they can process vast amounts of data quickly and efficiently, making them suitable for large-scale verification tasks. However, their accuracy can be questionable in many cases, ranging from struggles with nuanced or context-specific content (Boukouvalas and Shafer, 2024), issues with transfer and generalization (Huang et al., 2020; Pelrine et al., 2021, 2023), or just generically poor performance (e.g., even state-of-the-art methods often have below 70% accuracy compared to human labels (Zhang and Gao, 2023; Pelrine et al., 2023)). Thus, the quality of algorithmic labels is often dubious.
Appendix B Data Quality Supplement
B.1. Supplement on Keyword Analysis
Table 6 displays the results of the random forest classifier, including True and False (T-F) labels, as well as True, False and Mixed (T-F-M) labels for each dataset, alongside their corresponding baselines. Datasets located above the solid line represent the claim datasets and those positioned below the line correspond to the paragraph datasets.
Table 6. Keywords correlations evaluation. A high predictivity score which far exceeds its corresponding baseline, means that the keywords provide an unrealistically strong prediction. All numbers are % macro F1.
Dataset Keywords Predictivity T-F Baseline T-F Keywords Predictivity T-F-M Baseline T-F-M
Check-COVID 44.9 47.3 - -
ClaimsKG 44.0 50.1 27.0 33.6
Climate-Fever 44.9 50.2 26.7 33.4
CoAID 60.9 50.2 - -
COVID-19-Rumor 48.9 50.7 - -
COVID-FACT 40.6 52.3 - -
DeFaktS 37.1 49.9 - -
FakeCovid 49.9 49.9 32.7 32.8
FaVIQ 34.8 50.4 - -
FEVER 41.3 50.0 - -
FEVEROUS 37.4 49.5 - -
FibVID 53.3 49.6 - -
HoVer 38.1 49.5 - -
IFND 82.2 50.0 - -
LIAR 39.5 49.3 22.9 34.5
LIAR-New 48.1 50.0 31.4 31.9
MM-COVID 77.1 51.2 - -
NLP4IF-2021 48.8 52.4 - -
PubHealthTab 42.6 49.5 - -
Rumors 45.6 51.3 29.0 34.7
Snopes Fact-news 44.1 51.7 27.6 31.8
TruthSeeker2023 66.8 50.0 - -
Twitter15 62.2 48.9 - -
Twitter16 66.4 43.3 - -
Verite 39.9 49.0 - -
X-Fact 45.8 49.8 29.2 33.5
BanFakeNews 91.8
BenjaminPoliticalNews 72.2
Celebrity 64.0
CT-FAN 62.2
FA-KES 51.5
FakeNewsAMT 45.7
ISOT Fake News 91.8
TI-CNN 89.1
@@ B.1. Supplement on Keyword Analysis
Furthermore, figures  8,  10,  12 and  14 show the distribution of the 40 most frequent words across the datasets IFND, MM-COVID, Truthseeker2023, and Twitter16, the four claim datasets with the highest macro F1 score (%). The prevalence of the words in each veracity category was calculated using their relative frequency. A word positioned at x = 1 indicates that it is systematically associated with the veracity category specified by the label, meaning that 100% of statements containing this word are labeled the same way. Additionally, figures  810,  12, and  14, plotted using ScatterText, highlight the words that had a significant impact on the random classifier presented in Table  6, with color indicating the frequency of a word’s association with a label.
Table 7. Identification of spuriously predictive keywords.
@@
Figure 7. IFND
Figure 8. IFND Predictivity
Figure 9. MM-COVID
Figure 10. MM-COVID Predictivity
Figure 11. Truthseeker2023
Figure 12. Truthseeker2023 Predictivity
Figure 13. Twitter16
Figure 14. Twitter16 Predictivity
B.2. Supplement on Temporal Analysis
Table 8 presents the results of the spurious temporal correlations discussed in Section 3. Similar to the keywords analysis, the datasets located above the solid line include those with claims, whereas those below the line consist of paragraph datasets.
Table 8. Temporal correlations evaluation. A high score here means time—and information correlated with it—is unrealistically predictive.
Dataset Evaluation Type Temporal Predictivity (F1)
ClaimsKG Date 62.3
CoAID Date 48.3
DeFaktS Date 37.4
FakeCovid Date 49.9
FibVID Date 62.2
LIAR-New Date 53.7
Rumors Date 74.2
X-Fact Date 61.5
AntiVax TweetID 46.8
CMU-MisCOV19 TweetID 45.6
Covid-19-disinformation TweetID 46.5
MediaEval TweetID 72.2
Twitter15 TweetID 85.6
Twitter16 TweetID 95.9
WICO TweetID 40.7
BanFakeNews Dates 98.3
FA-KES Dates 50.1
IsotFakeNews Dates 87.7
TI-CNN Dates 78.1
B.3. Feasibility Evaluation
B.3.1. Claims Feasibility Evaluation
The prompt where it is explicitly indicated that the AI veracity assessment system has access to web search is as follows:
The following statement is going to be given to an AI system to determine if its true or false and write an explanation why.
Statement: ’{statement}’
The only thing the AI will be given is the statement itself, as written above - no context, visuals, or any other information. Your task is to assess if the AI could possibly give a valid answer. Note that this is not about assessing how likely the AI is to give the right answer, but whether its even possible to evaluate the veracity of the statement based on the information given. The AI will have access to a web search system to look for both primary and secondary sources, but the evaluation might still be impossible if there is too much ambiguity or missing context.
For example, here is a non-exhaustive list of information that might make it hard to evaluate the veracity of a statement if missing:
1. Identity of a key person, such as the speaker or someone else referenced ambiguously in the statement.
2: Location, if veracity depends on it but it isnt provided.
3. Textual information or evidence thats mentioned in the statement but not supplied.
4. Visual or audio evidence mentioned in the statement (note that the AI will only be given the statement text).
5. Temporal information. Note that the date the statement was made is unknown. This might not be relevant, though, if the statement could be evaluated as true or false regardless of when it was made.
6. Theres no claim for which evaluating the veracity even makes sense.
Rate on the following scale how possible it seems to evaluate the veracity of the statement:
1: Feasible, assuming that the retrieval of external knowledge is possible- There is some clear ambiguity, missing context, or multiple potential interpretations. But there seems to be around one-half chance of evaluating the meaning as intended or figuring out the context from a strong knowledge base or web search.
0: Impossible to evaluate, even with access to external knowledge retrieval systems. There are clearly multiple valid ways the statement could be interpreted that would strongly influence the veracity, mandatory and irrecoverable information is missing, or the statement contains no claim or is downright nonsensical.
Give a brief explanation, then write a vertical bar ”|”, followed by your rating as a number alone.
The “search disabled” version, where it is not specified explicitly that the AI has access to web search, is exactly the same as above except omitting the sentence “The AI will have access to a web search system to look for both primary and secondary sources, but the evaluation might still be impossible if there is too much ambiguity or missing context.”
We count a claim as AI-labeled “feasible” as long as the AI marked the claim as either feasible with search or as feasible- no search required.
B.3.2. Paragraphs Feasibility Evaluation
The prompt used for feasibility evaluation on paragraphs datasets is the same as the one for shorter claims.
B.3.3. Copy of Human Annotator Instructions
The concept of claim feasibility is inspired by point 4 (”Methodology”) of
https://arxiv.org/abs/2401.01197. As mentioned in this paper, resolving ambiguity in a statement can be facilitated through web retrieval, particularly when context is missing but can be inferred. For example, when an unknown person is mentioned in a specific event (e.g., ”politician X held a press conference about COVID on September 12, 2021 at the White House”).
In other words, the feasibility label indicates whether a language model (LLM) has enough information to assess the truthfulness of a claim. To achieve this, three categories are defined:
Feasible
The statement provides enough context for the LLM to determine its truthfulness with certainty.
Example 1: ”Bill Clinton death ruled a homicide, death by poison.”
The statement contains all necessary information: the person (Bill Clinton) and the event (his alleged death by poisoning). Sufficient information is present to verify the claim (and we know Bill Clinton is still alive).
Example 2: ”Ohio State scored fewer points than Purdue at the 1947 NCAA Swimming and Diving Championships.”
The statement includes sufficient details (institution, year, competition) to allow for factual verification without additional context.
Feasible with Web Search
Some key information is missing, preventing the LLM from determining the claim’s truthfulness without retrieving additional data online.
Example: ”A law allows people to go for a run during the state of alarm in Spain.”
The statement references a specific law, but it is not explicitly identified. A web search would be necessary to locate the relevant legal text and verify the claim. Since the country (Spain) and general content of the law are mentioned, this facilitates an easy online search.
Not Feasible
The statement is too vague or incomplete for a web search to provide sufficient evidence for verification.
Example: ”The (COVID-19) cases are going up but it’s because the testing is going up.”
The statement lacks crucial details such as the time period and location, making factual verification impossible. The claim might be false if based on misinformation disseminated during the COVID-19 pandemic, or it might be true if made during a period when increased testing corresponded to rapid virus spread.
B.3.4. Human Annotator Team and Aggregation
Our annotating team included 8 human experts: 3 authors and 5 colleagues of the authors. Each annotator completed an approximately equal number of examples (2̃90).
When aggregating these annotations, we note that this setting is slightly different from typical ones where the majority vote is expected to converge to the true label. Here, a claim can be proven infeasible by counterexample: demonstrating that there are two possible contexts it could refer to. This means that a minority of annotators who think of a counterexample could be correct, even while a majority misses the ambiguity.
This creates a challenge for determining how to combine annotations from multiple annotators. In our annotation process, to obtain both examples with multiple annotators and maximize how much of each dataset we could cover, approximately 25% was distributed to have a single annotator, 60% to have two annotators, and the remainder had three. In the aggregation process, we first convert data that was labeled as feasible with or without search to just “feasible”. The key question then is how to tiebreak cases where two annotators disagree.
We consider three options. First, we could tie-break in favor of “not feasible”, setting a lower bound on overall feasibility. This may be the correct measure, by the logic above. However, we could also set a generous upper bound by tie-breaking in favor of “feasible”. There does not seem to be clear reasoning supporting this upper bound being the true value, but it does provide a stress test for our arguments that feasibility is a significant issue. Finally, we could set a middle ground by taking the average of the two.
In Figure 15, we present each of these measures. We see that even with the most generous, upper bound assessment, many datasets still have a great deal of infeasible examples. Meanwhile, the lower bound suggests many datasets could have incredibly low feasibility, to an extent that predicting veracity on the text of these datasets is not only partially but almost entirely predicting noise. In the main paper, we take the moderate approach with the average. We note, though, that even more extensive investigation of this phenomenon could be a good area for future work, and might reveal an even more severe problem than we highlight in our main paper.
Figure 15. Comparison between human experts and automated evaluation of feasibility, aggregated based on 40 examples sampled from each data source for human annotations, and 300 for automated evaluation.
Appendix C Baselines and Metrics Supplement
C.1. Implementation Details of GPT-4 with Web Search Predictive System